sum of ages is 6 if the older ages 4 more than the younger how old are they

Answer by vheroli(126)

First, we will disregard their difference in age by subtracting their difference in age:
6-4=2
Then, divide it by 2 since you are comparing 2 people
2/2=1
Each one gets 1 year now:
Age of Older = 1
Age of Younger = 1
But it still not the final answer since we still need to apply their difference in age which we disregard earlier.
If you analyze, to get the age of the older you must add 4:
4+1=5
so
Age of Older = 5
Age of Younger = 1
It is now your final answer since it adds up to 6 and their difference is 4.